The base edge of a regular pentagonal prism is 8cm and its base are 16cm apart. Find its volume in cubic cm.
Answer by Alan3354(69443)

The base edge of a regular pentagonal prism is 8cm and its base are 16cm apart. Find its volume in cubic cm.
-------------
Vol = base area*h (h = 16 cm)
-----
Area = ns^2*cot(180/n)/4 where n = # of sides, s = side length
---
Vol = (5*64*cot(36)/4)*16 cc